OpenStudy (anonymous):

Solve each quadratic equation. Show your work Where do i start? 4x^2 = 100

OpenStudy (welshfella):

divide both sides by 4 to isolate x^2:- x^2 = 25 can you continue?

OpenStudy (anonymous):

that will give the answer right

OpenStudy (welshfella):

well you need to find the value of x x ^2 means x times x so you need to think of a number which when multiplied by itself gives 25 ?

OpenStudy (anonymous):

Okay. 4x^2/4 100/4 = 25

OpenStudy (anonymous):

Does that get rid of the x^2?

OpenStudy (anonymous):

the work shown should help get the solution which i think is 25

OpenStudy (anonymous):

Yeah 100/4=25

OpenStudy (anonymous):

Do i have x^2 any more or does it go with the 4 @welshfella ?

OpenStudy (welshfella):

no 25 is not the solution we have x^2 = 25 so x = ?

OpenStudy (anonymous):

so 25 is the answer

OpenStudy (anonymous):

25x^2?

OpenStudy (anonymous):

100 is the solution

OpenStudy (anonymous):

x=4

OpenStudy (welshfella):

no the solution is the value of x we know that x times x = 25 so what must x be?

OpenStudy (welshfella):

Not 4 because 4*4 = 16 not 25

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

How do we find the value of x?

OpenStudy (welshfella):

x is the square root of 25 that is the number that when multiplied by itself gives 25 for example 3 is square root of 9 as 3 * 3 = 9

OpenStudy (anonymous):

hold on let me try to figure this out

OpenStudy (welshfella):

what times what = 25?

OpenStudy (anonymous):

so you would use square root

OpenStudy (welshfella):

yea

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

625?

OpenStudy (anonymous):

i think so

OpenStudy (welshfella):

no - thats the SQUARE of 25 25*25 = 625

OpenStudy (anonymous):

okay what do you think

OpenStudy (welshfella):

5 *1 = 5 5*2 = 10 5*3 = 15 5*4 = 20 what comes next?

OpenStudy (welshfella):

now you have it

OpenStudy (anonymous):

25

OpenStudy (anonymous):

sooooo

OpenStudy (welshfella):

5 is the square root of 25 because 5 * 5 = 25 x^2 = 25 so x = 5

OpenStudy (anonymous):

Alright

OpenStudy (welshfella):

actually there is 2 answers 5 and -5 because -5*-5 is also equal to 25

OpenStudy (anonymous):

okay was i right

OpenStudy (anonymous):

Neat. Thanks for helping me i appreciate it :)

OpenStudy (welshfella):

@kitty_lover1234 No . 25 is not the answer

OpenStudy (anonymous):

what do you think it is

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

okay

OpenStudy (welshfella):

also -5 write the answer as {5,-5}

OpenStudy (anonymous):

Alright

OpenStudy (welshfella):

that is what is known as the 'solution set'. Written in curly brackets.

OpenStudy (anonymous):

i see where i went wrong